OpenStudy (anonymous):

solve the system of equations and give your answer as an ordered pair y-5x=-24 -4y-2x=8

OpenStudy (anonymous):

The given eq are y-5x=-24 ie y=5x-24---(1) -4y-2x=8 ------(2) from 1 & 2 by substitution method we have: -4(5x-24)-2x=8 i.e. -20x-2x+96 =8 i.e. -22x =8 -96 i.e. -22x =-88 i.e. x=-88/-22 i.e. x=4 substituting x=4 in eq i we find y=5x-24=5*4-24=20-24=-4 thus (x,y)= (4, -4) is the ordered pair solution of the given system of equation

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

medal pleasse